A Jar contains 24 red marbles numbered 1 to 24 and 50 blue marbles numbered 1 to 50. A marble is drawn at random from the Jar. Find the probability of the given event. Please enter reduced fractions.


a) The marble is red. P(red)=

b) The marble is odd numbered. P(odd)=

c) The marble is red or odd numbered. P(red or odd)=

d)The marble is blue or even numbered. P(blue or even)=


1
Expert's answer
2022-05-02T16:16:15-0400

a)

"P(Red)=\dfrac{24}{24+50}=\dfrac{12}{37}"

b)

"P(Odd)=\dfrac{12+25}{24+50}=\dfrac{1}{2}"

c)


"P(Red\cup Odd)=\dfrac{24+25}{24+50}=\dfrac{49}{74}"

d)


"P(Blue\cup Even)=\dfrac{12+50}{24+50}=\dfrac{31}{37}"